What discount rate would make you indifferent between receiving $3,492.00 per year forever and $5,668.00 per year for 29.00 years? Assume the first payment of both cash flow streams occurs in one year.

Also if there is a way to input it on calculator without long math, please include (with keystrokes). Thanks
Answer #1

This can be solved as under

PV of perpetuity = PV of Annuity

A/r = A[1-1/(1+r)^n / r]

A = annuity, r = rate of interest , n = number of period

3492/r = 5668[1-1/(1+r)^29 /r]

3492/5668 = r[[1-1/(1+r)^29 /r]

0.61609 =1-1/(1+r)^29

1/(1+r)^29 = 1-0.61609

1/(1+r)^29=0.3839

(1+r)^29 = 2.60477

(1+r) = 2.60477^1/29

(1+r) = 1.03356

r = 0.03356

ie 3.356%